OpenStudy (anonymous):

Solve the system using the substitution method. If there is no solution, state so. Can someone please help me to understand how to do this? y=2x+5 x+y=2

OpenStudy (anonymous):

where is equation

OpenStudy (anonymous):

Sure but I need an example to help work with. Do you have one?

OpenStudy (anonymous):

I have one hold on.

OpenStudy (anonymous):

y=2x+5,so x+2x+5=2,now solve

OpenStudy (anonymous):

y=2x+5 and x+y=2

OpenStudy (anonymous):

OH this is fine Now to substitute means to replace one thing with an another...So what are we substituting.? Well that would be an amount for a variable. To solve an equation you can really only have one variable. y=2x +5 right? ...so replace y with 2x +5 in the other equation so you have only one variable. x + 2x + 5 = 2 3x + 5 = 2 Subtract 5 and you get -3 3x = -3 Now divide by 3 and you get x = -1 Sorry it was so many lines..but i hope this helps

OpenStudy (anonymous):

oh ok @Mpost1994 @roshan2004 Now I see what I was doing wrong I was trying to solve for x but didn't know I could also use y to try and solve the problem explains why it was working lol This is so confusing!

OpenStudy (anonymous):

How do you know which variable to use?

OpenStudy (anonymous):

You can use any variable you like, you will get the same answer, Don't worry !!!, that's the beauty of equation :-)

OpenStudy (anonymous):

You can use any variable you like.. but most of the time it will be obvious which one to use...for example I knew to use y this time because it told me exactly what y = in terms of x that is ..what y equals using only numbers and the x variable. Do you get it? :)

OpenStudy (anonymous):

Yes but is this a solution? @Mpost1994 @roshan2004 Because on my hw I am suppose to write either solution-independent no solution-parallel or same line infinitely many dependent how do I know which one to put for this system?

OpenStudy (anonymous):

Or how can I determine the answer?

OpenStudy (anonymous):

Can you give me the choices seperately again?

OpenStudy (anonymous):

A. solution-independent B. no solution-parallel C. same line infinitely many dependent

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Don't you need to solve for y roshan?

OpenStudy (anonymous):

@roshan2004 Why A? How did you get that? Can you please explain?

OpenStudy (anonymous):

If it's a solution wouldn't there be and x and y-intercepts?

OpenStudy (anonymous):

Look at this site and maybe this will explain...alright? http://www.algebra.com/algebra/homework/coordinate/Types-of-systems-inconsistent-dependent-independent.lesson Pm after you read this okay?

OpenStudy (anonymous):

Here's an example I have that the professor did on the board: Solve by substitution : 3x+y=-1 x+2y=8 1. x+2y=8 -2y -2y x=-2y+8 2. 3(-2y+8)+y=-1 -6y+24+y=-1 -5y+24=-1 -24 -24 -5y=-25 / -5 y=5 3. 3x+5=-1 -5 -5 3x=-6 x=-2

OpenStudy (anonymous):

So solution is (-2,5) Independent But idk how he did this?

OpenStudy (anonymous):

Oh well that would what we already did .....just solve for both variables using both equations

OpenStudy (anonymous):

y= -x+2 y= 2x + 5 Put those into a graph and solve for the solution ..which is where the lines cross

OpenStudy (anonymous):

so y=3?

OpenStudy (anonymous):

@Mpost1994

OpenStudy (anonymous):

Yes :) -(-1) +2 or 2(-1) +5 both come out to 3 ^.^